# Towards Real World Debiasing: A Fine-grained Analysis On Spurious Correlation

### v2

@@ 5.4 Analysis
As we do not know how biased or is the training data biased at all in real-world scenarios, it is important to evaluate the performance of debias methods on unbiased training data to ensure that they do not cause severe performance degradation, if not improving the performance. As shown in Table 3, existing DBAM methods perform poorly on unbiased training data, causing severe performance degradation, yet our approach greatly boosts their performances.
Hyper-parameter sensitivity.
We examine three feature destruction methods: pixel-shuffling, patch-shuffling, and center occlusion, to destroy object shapes. We observed that patch-shuffle with patch-size 8 exhibits the best performance on Corrupted CIFAR10 which is of size 32x32.
@@
Algorithm Colored MNIST Corrupted CIFAR10 ERM 94.14 ± 0.21 67.91 ± 0.13 LfF 70.19 ± 1.50 52.04 ± 2.14 + Ours 93.18 ± 0.26 57.29 ± 0.22 LfF BE 84.14 ± 0.61 55.64 ± 0.66 + Ours 90.02 ± 0.54 56.28 ± 0.45 DisEnt 75.24 ± 3.40 58.50 ± 0.20 + Ours 92.24 ± 0.44 64.58 ± 0.02 DisEnt BE 80.66 ± 0.90 58.57 ± 0.12 + Ours 89.10 ± 1.28 62.97 ± 0.16
$T_{fd}$ param BC Avg N/A N/A 47.70 ± 3.58 54.15 ± 3.02 pixel-shuffle 1 51.44 ± 1.01 55.43 ± 0.20 patch-shuffle 2 51.07 ± 0.48 55.29 ± 0.27 4 49.41 ± 0.26 55.40 ± 0.26 8 54.81 ± 0.74 63.06 ± 0.77 16 49.74 ± 1.10 53.69 ± 0.31 center-occlusion 8 45.19 ± 1.41 51.61 ± 1.31 16 47.26 ± 0.54 50.94 ± 0.59 24 49.00 ± 0.80 52.60 ± 0.55 32 52.44 ± 0.87 55.76 ± 0.16
Figure 5: The performance of debias methods under various bias magnitudes and prevalence.
@@ 6 Conclusions and Discussion
In this work, we emphasize the importance of debias within the real world. To tackle real-world biases, we first proposed a fine-grained analysis framework to analyze dataset biases, based on which we further proposed a systematic evaluation framework for benchmarking debias methods under real-world biases. According to our result, we identified the insufficiency of existing methods and proposed a new approach to resolve it. Our experiments demonstrate the effectiveness of our approach. In Appendix G, we further discuss the limitations and future directions of this work.
@@ Appendix A More visualizations of biased distributions
CelebA [36] is a dataset for face recognition where each sample is labeled with 40 attributes, which has been adopted as a benchmark for debias methods. Following the experiment configuration suggested by Nam et al. [32], we focus on HeavyMakeup attributes that are spuriously correlated with Gender attributes, i.e., most of the CelebA images with heavy makeup are women. As a result, the biased model suffers from performance degradation when predicting males with heavy makeup and females without heavy makeup. Therefore, we use Heavy_Makeup as the target attribute and Male as a spurious attribute. The joint distribution between the Male and Heavy_Makeup attribute of the CelebA dataset is plotted in Figure 6(a). It is clear that the biased distribution of CelebbA aligns with that in other existing benchmarks, forming a "diagonal distribution".
WaterBirds.
WaterBirds [16] is a synthetic dataset with the task of classify images of birds as "waterbird" and "landbird", which is adopted as a benchmark for debias methods. The label of WaterBirds is spuriously correlated with the image background, i.e. Place attribute, which is either "land" or "water". The joint distribution between the Place and Bird attribute of the WaterBirds dataset is plotted in Figure 6(b).
@@
The Adult [37] dataset, also known as the "Census Income" dataset, is widely used for tasks such as income prediction and fairness analysis. Each sample is labeled with demographic and income-related attributes. The dataset has been adopted as a benchmark for debias methods, particularly focusing on the correlation between race and income. The joint distribution between Race and Income attributes of the Adult dataset is plotted in Figure 6(c). It is clear that the biased distribution of Adult does not align with that of other existing benchmarks.
@@
The German [38] dataset, also known as the "German Credit" dataset, is commonly used for credit risk analysis and fairness studies. Each sample is labeled with various attributes related to creditworthiness. The dataset serves as a benchmark for debias methods, emphasizing the correlation between age and creditworthiness. The joint distribution between Age and Creditworthiness attributes of the German dataset is plotted in Figure 6(d). It is clear that the biased distribution of German does not align with that of other existing benchmarks.
(a) CelebA
(b) WaterBirds
@@
Figure 6: Visualization of the joint distribution for datasets, where the y-axis is the target attribute and the x-axis is the spurious attribute. Figure 6(a) and 6(b) visualize the distribution of existing benchmarks. Figure 6(c) and 6(d) visualize the distribution of real-world datasets. The biased distribution of existing benchmarks and real-world datasets is not alike.
@@ D.3 Baselines
LfF.
Learning from Failure (LfF) [12] is a debiasing technique that addresses the issue of models learning from spurious correlations present in biased datasets. The method involves training two neural networks: one biased network that amplifies the bias by focusing on easily learnable spurious correlations, and one debiased network that emphasizes samples the biased network misclassifies. This dual-training scheme enables the debiased network to focus on more meaningful features that generalize better across various datasets.
DisEnt.
The DisEnt [9] method enhances debiasing by using disentangled feature augmentation. It identifies intrinsic and spurious attributes within data and generates new samples by swapping these attributes among the training data. This approach significantly diversifies the training set with bias-conflicting samples, which are crucial for effective debiasing. By training models with these augmented samples, DisEnt achieves better generalization and robustness against biases in various datasets.
BE.
BiasEnsemble (BE) [20] is a recent advancement in debiasing techniques that emphasizes the importance of amplifying biases to improve the training of debiased models. BE involves pretraining multiple biased models with different initializations to capture diverse visual attributes associated with biases. By filtering out bias-conflicting samples using these pre-trained models, BE constructs a refined bias-amplified dataset for training the biased network. This method ensures the biased model is highly focused on bias attributes, thereby enhancing the overall debiasing performance of the subsequent debiased model.
@@ D.4 Implementation details
Following [12, 9], we use a multi-layer perceptron (MLP) which consists of three hidden layers for Colored MNIST. For the Corrupted CIFAR10 dataset, we train ResNet18 [42] with random initialization.
@@
D.5 Applying DiD to DBAM methods
@@ Appendix E Additional Empirical Results
E.1 Detailed results of the main experiments
@@
Distr. Algorithm Accuracy BA acc BC acc BN acc Avg acc LMLP ERM 80.40 ± 0.81 62.50 ± 0.15 - 64.29 ± 0.06 LfF 59.13 ± 0.68 55.03 ± 0.04 - 55.44 ± 0.09 + DiD 69.47 ± 0.96 62.04 ± 0.21 - 62.78 ± 0.10 LfF BE 70.87 ± 1.30 52.10 ± 0.30 - 53.98 ± 0.40 + DiD 63.23 ± 2.10 53.21 ± 0.20 - 54.21 ± 0.38 DisEnt 61.58 ± 0.57 55.45 ± 0.23 - 56.06 ± 0.17 + DiD 72.23 ± 0.74 60.84 ± 0.40 - 61.98 ± 0.30 DisEnt BE 62.73 ± 0.61 56.59 ± 0.08 - 57.20 ± 0.13 + DiD 65.98 ± 0.40 60.92 ± 0.20 - 61.42 ± 0.21 HMLP ERM 84.67 ± 0.64 55.85 ± 0.17 65.75 ± 0.00 65.05 ± 0.13 LfF 73.33 ± 1.67 47.70 ± 0.58 54.58 ± 0.49 54.15 ± 0.41 + DiD 78.67 ± 2.14 54.81 ± 2.26 63.71 ± 2.69 63.06 ± 2.63 LfF BE 70.33 ± 2.19 50.96 ± 2.35 54.14 ± 0.25 54.02 ± 0.36 + DiD 68.80 ± 0.88 50.20 ± 0.79 54.39 ± 0.18 54.15 ± 0.15 DisEnt 61.67 ± 1.67 52.48 ± 0.56 54.65 ± 0.56 54.53 ± 0.49 + DiD 73.67 ± 2.64 55.26 ± 0.93 62.11 ± 0.17 61.61 ± 0.13 DisEnt BE 75.33 ± 5.21 49.15 ± 1.54 56.86 ± 0.30 56.35 ± 0.35 + DiD 78.40 ± 1.00 54.09 ± 1.07 62.05 ± 0.34 61.50 ± 0.38 HMHP ERM 89.97 ± 0.34 29.37 ± 0.30 - 35.43 ± 0.24 LfF 72.70 ± 0.81 35.30 ± 0.33 - 39.04 ± 0.33 + DiD 82.07 ± 1.09 37.05 ± 0.31 - 41.55 ± 0.19 LfF BE 82.73 ± 0.92 31.48 ± 0.82 - 36.61 ± 0.65 + DiD 78.30 ± 0.47 32.90 ± 1.79 - 37.44 ± 1.61 DisEnt 70.77 ± 2.27 36.04 ± 0.62 - 39.51 ± 0.36 + DiD 76.60 ± 0.70 39.05 ± 0.35 - 42.80 ± 0.25 DisEnt BE 78.60 ± 1.56 34.20 ± 0.43 - 38.64 ± 0.38 + DiD 78.70 ± 1.47 37.72 ± 0.96 - 41.82 ± 0.91

### v3

@@ 5.3 Analysis
As we do not know how biased or is the training data biased at all in real-world scenarios, it is important to evaluate the performance of debiasing methods on unbiased training data to ensure that they do not cause severe performance degradation, if not improving the performance. As shown in Table 4, existing DBAM methods perform poorly on unbiased training data, causing severe performance degradation, yet our approach greatly boosts their performances.
@@
Table 3: Results on 3 datasets with more complex and realistic sets of features further show the effectiveness of our approach.
Algorithm BAR NICO WaterBirds ERM 35.32 ±0.27 42.61 ±0.33 56.53 ±0.27 LfF 37.73 ±1.00 51.69 ±3.06 50.02 ±0.00 + Ours +3.34 ±1.69 +2.80 ±3.10 +4.47 ±0.13 DisEnt 59.11 ±1.75 39.73 ±0.58 56.75 ±4.19 + Ours +3.92 ±0.62 +16.55 ±1.29 +11.29 ±0.69 BE LfF 38.40 ±0.65 44.09 ±1.95 52.98 ±0.28 + Ours +1.08 ±1.67 +8.23 ±1.49 +1.92 ±0.12 BE DisEnt 62.74 ±1.23 39.58 ±0.91 53.85 ±2.14 + Ours +0.70 ±1.20 +13.50 ±1.62 +1.99 ±3.65
@@
Algorithm Colored MNIST Corrupted CIFAR10 ERM 94.14 ± 0.21 67.91 ± 0.13 LfF 70.19 ± 1.50 52.04 ± 2.14 + Ours 93.18 ± 0.26 57.29 ± 0.22 DisEnt 75.24 ± 3.40 58.50 ± 0.20 + Ours 92.24 ± 0.44 64.58 ± 0.02 BE LfF 84.14 ± 0.61 55.64 ± 0.66 + Ours 90.02 ± 0.54 56.28 ± 0.45 BE DisEnt 80.66 ± 0.90 58.57 ± 0.12 + Ours 89.10 ± 1.28 62.97 ± 0.16
Figure 5: The performance of debiasing methods under various bias magnitudes and prevalence.
@@ 6 Conclusions and Discussion
In this work, we emphasize the importance of debias within the real world. To tackle real-world biases, we first proposed a fine-grained analysis framework to analyze dataset biases, based on which we further proposed a systematic evaluation framework for benchmarking debiasing methods under real-world biases. According to our result, we identified the insufficiency of existing methods and proposed a new approach to resolve it. Our experiments demonstrate the effectiveness of our approach. In Appendix G, we further discuss the limitations and future directions of this work.
@@ Appendix A More visualizations of biased distributions
WaterBirds.
WaterBirds [16] is a synthetic dataset with the task of classify images of birds as "waterbird" and "landbird", which is adopted as a benchmark for debiasing methods. The label of WaterBirds is spuriously correlated with the image background, i.e. Place attribute, which is either "land" or "water". The joint distribution between the Place and Bird attribute of the WaterBirds dataset is plotted in Figure 6(a).
@@
CelebA [29] is a dataset for face recognition where each sample is labeled with 40 attributes, which has been adopted as a benchmark for debiasing methods. Following the experiment configuration suggested by Nam et al. [32], we focus on HeavyMakeup attributes that are spuriously correlated with Gender attributes, i.e., most of the CelebA images with heavy makeup are women. As a result, the biased model suffers from performance degradation when predicting males with heavy makeup and females without heavy makeup. Therefore, we use Heavy_Makeup as the target attribute and Male as a spurious attribute. The joint distribution between the Male and Heavy_Makeup attribute of the CelebA dataset is plotted in Figure 6(b). It is clear that the biased distribution of CelebbA aligns with that in other existing benchmarks, forming a "diagonal distribution".
@@
The Adult [39] dataset, also known as the "Census Income" dataset, is widely used for tasks such as income prediction and fairness analysis. Each sample is labeled with demographic and income-related attributes. The dataset has been adopted as a benchmark for debiasing methods, particularly focusing on the correlation between race and income. The joint distribution between Race and Income attributes of the Adult dataset is plotted in Figure 6(c). It is clear that the biased distribution of Adult does not align with that of other existing benchmarks.
@@
The German [40] dataset, also known as the "German Credit" dataset, is commonly used for credit risk analysis and fairness studies. Each sample is labeled with various attributes related to creditworthiness. The dataset serves as a benchmark for debiasing methods, emphasizing the correlation between age and creditworthiness. The joint distribution between Age and Creditworthiness attributes of the German dataset is plotted in Figure 6(d). It is clear that the biased distribution of German does not align with that of other existing benchmarks.
The description of MultiNLI and CivilComments-WILDS can be found in Appendix E.
(a) WaterBirds
(b) CelebA
@@
(e) MultiNLI
(f) CivilComments-WILDS
Figure 6: Visualization of the joint distribution for datasets, where the y-axis is the target attribute and the x-axis is the spurious attribute. Figure 6(a) visualize the distribution of existing benchmarks. Figure 6(b), 6(c), 6(d), 6(e), and 6(f) visualize the distribution of real-world datasets. The biased distribution of existing benchmarks and real-world datasets is not alike.
@@ D.2 Datasets
BAR [12].
Biased Action Recognition (BAR) is a semi-synthetic dataset deliberately curated to contain spurious correlations between six human action classes and six place attributes. Following Nam et al. [12], the ratio of bias-conflicting samples in the training set was set to 5%, and the test set consisted of only bias-conflicting samples. We report the accuracy of bias-conflicting samples following [12].
NICO [15]
NICO is a real-world dataset for simulating out-of-distribution image classification scenarios. Following the setting used by Wang et al. [44], we use a curated animal subset of NICO that exhibits strong biases (thus still semi-synthetic), which is labeled with 10 object and 10 context classes for evaluating the debiasing methods. The training set consists of 7 context classes per object class and they are long-tailed distributed (e.g., dog images are more frequently coupled with the ‘on grass’ context than any of the other 6 contexts). The validation and test sets consist of 7 seen context classes and 3 unseen context classes per object class. We verify the ability of debiasing a model from object-context correlations through evaluation on NICO. We report the average accuracy on the test set following [15].
WaterBirds [38].
The task is to classify images of birds as “waterbird” or “landbird”, and the label is spuriously correlated with the image background, which is either “land” or “water”. We report the worst group accuracy following [16].
MultiNLI [30].
Given a pair of sentences, the task is to classify whether the second sentence is entailed by, neutral with, or contradicts the first sentence. We use the spurious attribute from [38], which is the presence of negation words in the second sentence; due to the artifacts from the data collection process, contradiction examples often include negation words.
CivilComments-WILDS [3].
The task is to classify whether an online comment is toxic or non-toxic, and the label is spuriously correlated with mentions of certain demographic identities (male, female, White, Black, LGBTQ, Muslim, Christian, and other religion). We use the evaluation metric from [3], which defines 16 overlapping groups (a, toxic) and (a, non-toxic) for each of the above 8 demographic identities a, and report the worst-group performance over these groups.
@@ D.3 Baselines
LfF [12].
Learning from Failure (LfF) is a debiasing technique that addresses the issue of models learning from spurious correlations present in biased datasets. The method involves training two neural networks: one biased network that amplifies the bias by focusing on easily learnable spurious correlations, and one debiased network that emphasizes samples the biased network misclassifies. This dual-training scheme enables the debiased network to focus on more meaningful features that generalize better across various datasets.
DisEnt [9] .
The DisEnt method enhances debiasing by using disentangled feature augmentation. It identifies intrinsic and spurious attributes within data and generates new samples by swapping these attributes among the training data. This approach significantly diversifies the training set with bias-conflicting samples, which are crucial for effective debiasing. By training models with these augmented samples, DisEnt achieves better generalization and robustness against biases in various datasets.
BE [20].
BiasEnsemble (BE) is a recent advancement in debiasing techniques that emphasizes the importance of amplifying biases to improve the training of debiased models. BE involves pretraining multiple biased models with different initializations to capture diverse visual attributes associated with biases. By filtering out bias-conflicting samples using these pre-trained models, BE constructs a refined bias-amplified dataset for training the biased network. This method ensures the biased model is highly focused on bias attributes, thereby enhancing the overall debiasing performance of the subsequent debiased model.
JTT [16].
JTT is a classic DBAM method that has been applied to both the image and NLP domains. JTT identifies challenging examples by training an initial model using standard empirical risk minimization (ERM) and collecting misclassified examples into an error set. The second stage involves re-training the model while upweighting the error set to prioritize examples that the first-stage model struggled with. This approach aims to address performance disparities caused by spurious correlations, leading to better generalization across groups with minimal additional annotation costs.
Group DRO [38].
Group DRO is a supervised debiasing method aiming to improve the worst group accuracy. It is commonly used as an upper bound in the worst group accuracy for unsupervised methods.
@@ D.4 Implementation details
Following [12, 9], we use a multi-layer perceptron (MLP) which consists of three hidden layers for Colored MNIST. For the Corrupted CIFAR10, BAR, NICO, WaterBirds dataset, we train ResNet18 [45] with random initialization. For CelebA dataset, we train ResNet50 with random initialization, following [16]. For MultiNLI and CivilComments-WILDS datasets, we use Bert for training, following [16].
@@
D.5 Design of feature destructing methods
For in visual recognition tasks, the shape of objects is a basic element of human visual perception [36]. Therefore, the patch-shuffle destruction of shape [37] when capturing bias from visual recognition datasets is a feasible approach. We adopt the patch-shuffle approach for all the visual dataset within the paper except for CelebA. We apply a gray-scale transformation for CelebA as its recognition task is hair color. Anyhow, the feature destruction method could be highly flexible for different tasks.
For NLP tasks, we first introduce the common biases within the NLP domain followed by a simple design of feature destruction method in the NLP domain. The commonly used NLP datasets for debiasing are MultiNLI and CivilComments-WILDS dataset. Specifically, the bias within the MultiNLI dataset is the correlation between the negation words and the entailment task and the bias within the CivilComments-WILDS dataset is the correlation between words implying demographic identities and the toxicity task. The target features of both datasets are semantic information of the sentences where the position of words matters, and the spurious features are the individual words which is insensitive to positions. Furthermore, such position sensitivity difference between target and spurious features within NLP biases is not limited to these two datasets but rather quite common. For example, CLIP has also been found with the "bag of words" phenomenon [46], which ignores the semantic meaning of the inputs and relies on words individually for prediction. As a result, a straightforward approach for feature destruction is to shuffle the words within the sentences.
D.6 Applying DiD to DBAM methods
@@ Appendix E Additional Empirical Results
E.1 Detailed results and explanations of the main experiments
@@
The performance generally varies between different datasets, different types of biased distribution, and algorithms with and without BiasEnsemble, e.g. between LfF and BE LfF. Firstly, the inconsistency between datasets is likely to depend on how thoroughly the target feature is destroyed within the dataset. The target features of Colored MNIST, i.e. digits, are destroyed more completely by patch shuffling, for shape is the only feature within digits. In comparison, the target feature of Corrupted CIFAR10 is more complicated (including shape, texture, color, etc.), and thus can not be thoroughly destroyed by patch shuffling, causing relatively lower performance gain. Secondly, the performance inconsistency between different biased distributions is due to the reliance of existing DBAM methods on the high bias prevalence assumption for bias capturing as discussed in section 4.2. Specifically, as the bias prevalence of the training distribution becomes higher, better bias capture can be achieved by existing DBAM even without our method, thus making our improvement on the performance less significant. This conclusion is supported by our experimental results shown in Figure 5. As for the performance inconsistency between algorithms with and without BiasEnsemble, it is due to the fact that BiasEnsemble is also a method targeted to enhance the bias capture procedure of the debiasing framework. As we can see that BiasEnsemble is much more robust to the change in the bias magnitude and prevalence from Table 1. In other words, certain overlap between the goals of BiasEnsemble and our method resulted in smaller improvement of our method on BiasEnsemble-based baselines.
@@
Distr. Algorithm Accuracy BA acc BC acc BN acc Avg acc LMLP ERM 80.40 ± 0.81 62.50 ± 0.15 - 64.29 ± 0.06 LfF 59.13 ± 0.68 55.03 ± 0.04 - 55.44 ± 0.09 + DiD 69.47 ± 0.96 62.04 ± 0.21 - 62.78 ± 0.10 BE LfF 70.87 ± 1.30 52.10 ± 0.30 - 53.98 ± 0.40 + DiD 63.23 ± 2.10 53.21 ± 0.20 - 54.21 ± 0.38 DisEnt 61.58 ± 0.57 55.45 ± 0.23 - 56.06 ± 0.17 + DiD 72.23 ± 0.74 60.84 ± 0.40 - 61.98 ± 0.30 BE DisEnt 62.73 ± 0.61 56.59 ± 0.08 - 57.20 ± 0.13 + DiD 65.98 ± 0.40 60.92 ± 0.20 - 61.42 ± 0.21 HMLP ERM 84.67 ± 0.64 55.85 ± 0.17 65.75 ± 0.00 65.05 ± 0.13 LfF 73.33 ± 1.67 47.70 ± 0.58 54.58 ± 0.49 54.15 ± 0.41 + DiD 78.67 ± 2.14 54.81 ± 2.26 63.71 ± 2.69 63.06 ± 2.63 BE LfF 70.33 ± 2.19 50.96 ± 2.35 54.14 ± 0.25 54.02 ± 0.36 + DiD 68.80 ± 0.88 50.20 ± 0.79 54.39 ± 0.18 54.15 ± 0.15 DisEnt 61.67 ± 1.67 52.48 ± 0.56 54.65 ± 0.56 54.53 ± 0.49 + DiD 73.67 ± 2.64 55.26 ± 0.93 62.11 ± 0.17 61.61 ± 0.13 BE DisEnt 75.33 ± 5.21 49.15 ± 1.54 56.86 ± 0.30 56.35 ± 0.35 + DiD 78.40 ± 1.00 54.09 ± 1.07 62.05 ± 0.34 61.50 ± 0.38 HMHP ERM 89.97 ± 0.34 29.37 ± 0.30 - 35.43 ± 0.24 LfF 72.70 ± 0.81 35.30 ± 0.33 - 39.04 ± 0.33 + DiD 82.07 ± 1.09 37.05 ± 0.31 - 41.55 ± 0.19 BE LfF 82.73 ± 0.92 31.48 ± 0.82 - 36.61 ± 0.65 + DiD 78.30 ± 0.47 32.90 ± 1.79 - 37.44 ± 1.61 DisEnt 70.77 ± 2.27 36.04 ± 0.62 - 39.51 ± 0.36 + DiD 76.60 ± 0.70 39.05 ± 0.35 - 42.80 ± 0.25 BE DisEnt 78.60 ± 1.56 34.20 ± 0.43 - 38.64 ± 0.38 + DiD 78.70 ± 1.47 37.72 ± 0.96 - 41.82 ± 0.91
E.2 Hyper-parameter sensitivity
As shown in Table 9, we examine three feature destruction methods: pixel-shuffling, patch-shuffling, and center occlusion, to destroy object shapes. We observed that patch-shuffle with patch-size 8 exhibits the best performance on Corrupted CIFAR10 which is of size 32x32.
E.3 Application of DiD on the bias detection task
Some recent work [32, 47] has focused on the task of detecting biases rather than debiasing directly. Such methods also involve a biased auxiliary model for the detection. To test the effectiveness of DiD on bias detection tasks, we apply DiD to the recently proposed B2T [47] method. Specifically, B2T detects bias keywords by calculating their CLIP score, whose calculation involves a biased auxiliary model to define an error dataset, similar to JTT. A keyword is identified as biased if it has a higher CLIP score and the subgroup defined by it should have lower accuracy.
Following [47], we use CelebA as the dataset for bias detection, where the keyword "Actor" (a proxy for Male) is considered ground truth for class Blond, and the keyword "Actress" (a proxy for Female) is considered ground truth for the class not Blond. As we can see in Table 8, by applying DiD to the training of the auxiliary model, we effectively improve both metrics CLIP score and subgroup accuracy, enhancing B2T’s bias detection ability.
To further validate the effectiveness of DiD in improving the quality of the error dataset, we adopt the worst group precision and recall metrics proposed by [16] for evaluation. Specifically, the worst group precision and recall indicate how accurately the error dataset represents the worst group samples. As shown in Figure 8, DiD improves both worst group precision and recall, demonstrating better bias identification ability.
Table 8: DiD effectively improves the bias identification ability of B2T, improving both CLIP Score and Subgroup Accuracy on the ground truth bias keywords of CelebA dataset.
Blond: Actor Not Blond: Actress
[…]
B2T 0.125 86.71 2.188 97.11
B2T + DiD 0.188 85.29 2.297 95.81
(a) Worst group precision
(b) Worst group recall
Figure 8: DiD consistently improves the worst group precision and recall in the error dataset across the epochs.
$T_{fd}$ param BC Avg
N/A N/A 47.70 ± 3.58 54.15 ± 3.02
pixel-shuffle 1 51.44 ± 1.01 55.43 ± 0.20
patch-shuffle 2 51.07 ± 0.48 55.29 ± 0.27
4 49.41 ± 0.26 55.40 ± 0.26
8 54.81 ± 0.74 63.06 ± 0.77
16 49.74 ± 1.10 53.69 ± 0.31
center-occlusion 8 45.19 ± 1.41 51.61 ± 1.31
16 47.26 ± 0.54 50.94 ± 0.59
24 49.00 ± 0.80 52.60 ± 0.55
32 52.44 ± 0.87 55.76 ± 0.16
E.4 Results of BN samples under LMLP settings
To further examine the correctness of our analysis and the effectiveness of our design, we show the weights of BN samples under the LMLP settings. As the LMLP distribution defined in the main paper contains biased features with similar levels of bias magnitude, the choice of threshold for identifying BN samples becomes not so intuitive. Thus a threshold of 0 is selected for the categorization in the main paper, defining all samples either BA or BC samples. Consequently, we define another version of LMLP distribution named LMLP’ where the magnitude of bias for each feature is low but at the same time distinguishable from each other. (Please refer to Appendix B for details) Based on LMLP’ we are able to confidently define BN samples for the BN weights analysis. As shown in Figure 9, DiD consistently emphasizes BN samples in the LMLP distribution across datasets and debiasing algorithms.
(a) Colored MNIST
(b) Corrupted CIFAR10
Figure 9: DiD consistently emphasizes BN samples in LMLP distributions across datasets and algorithms. Our approach is marked with solid lines.
E.5 Addtional results on the WaterBirds dataset
As mentioned in Appendix D, the evaluations on the WaterBirds dataset are based on the ResNet18 architecture, which is the architecture widely adopted by many previous works [12, 9, 20]. However, there are also some other works [16] that evaluate the WaterBirds dataset based on the ResNet50 architecture with better baseline performances. To demonstrate that our approach is consistently effective regardless of the experimental settings, we further test our approach with the exact same setting in [16]. As shown in Table 10, DiD is consistently effective regardless of different experimental settings of WaterBirds.
Table 10: Results demonstrate that DiD is consistently effective regardless of different experimental settings of WaterBirds. The results are based on the ResNet50 architecture.
Bias supervision WaterBirds
Avg Acc. Worst-group Acc.
ERM No 78.82 31
JTT No 90.99 65.26
+DiD No +3.45 +17.45
Group DRO Yes 92.89 83.49